# Star formation beyond galaxies: widespread in-situ formation of intra-cluster stars in TNG50

### v1

@@ Abstract
We study the fraction of the intra-cluster light (ICL) formed in-situ in the three most massive clusters of the TNG50 simulation, with virial masses . We find that a significant fraction of ICL stars (-) are born in-situ. This amounts to a total stellar mass comparable to the central galaxy itself. Contrary to simple expectations, only a sub-dominant fraction of these in-situ ICL stars are born in the central regions and later re-distributed to more energetic orbits during mergers. Instead, many in-situ ICL stars form directly hundreds of kiloparsecs away from the central galaxy, in clouds condensing out of the circum-cluster medium. The simulations predict a present-date diffuse star formation rate of 1 /yr, with higher rates at higher redshifts. The diffuse star forming component of the ICL is filamentary in nature, extends for hundreds of kiloparsecs and traces the distribution of neutral gas in the cluster host halo. We discuss briefly how numerical details of the baryonic treatment in the simulation may play a role in this result and conclude that a sensitivity of erg s cm arcsec in H flux (beyond current observational capabilities) would be necessary to detect this diffuse star-forming component in galaxy clusters.
@@
The simplest models of galaxy formation suggest that most stars in a cluster should form in the high-density regions of individual galaxies. These models imply that the majority of stars in the intra-cluster light (ICL) originate from the tidal removal of stars from luminous galaxies with mass similar to the Milky Way (MW) (Puchwein et al., 2010; Cui et al., 2014; Cooper et al., 2015; Contini et al., 2013, 2018, 2019; Montenegro-Taborda et al., 2023), with relatively small contributions from dwarf mass galaxies, particularly in group environments with halo virial masses (Ahvazi et al., 2023). Nevertheless, observational studies (Barfety et al., 2022; Hlavacek-Larrondo et al., 2020) and some numerical simulations (Puchwein et al., 2010) indicate that a non-negligible fraction of ICL stars are formed in-situ.
In-situ ICL stars could have formed in the brightest cluster galaxy (BCG) galaxy and been later pushed onto highly energetic orbits by mergers. An alternative, more intriguing, possibility is that stars form directly in the circum-cluster medium (CCM), far away from the high density regions of the BCG. This low-density mode of star formation is potentially interesting, as it might occur under very different physical conditions than the normal star formation in galaxies, with consequences for the multi-phase nature of the CCM and its metallicity, as well as for the age distribution of ICL stars, especially if in-situ star formation continues to the present day. From an observational perspective, if a substantial amount of stars form outside galaxies, this would have strong implications for wide-field surveys in the UV or H, which are especially sensitive to the presence of young stars.
@@
In-situ star formation in the CCM is also predicted to be widespread at high-redshift, where AGN feedback might be less powerful and unable to offset the short timescales for cooling in the dense regions of the CCM (Webb et al., 2015b). In such a scenario, spontaneous run-away cooling in the CCM might lead to an increase of density high enough to trigger violent star-formation events in locations outside galaxies (Hlavacek-Larrondo et al., 2020).
@@
Interestingly, observations of off-center star formation at large distances from the BCG are consistent with this scenario for a few high- clusters (Webb et al., 2015a; Hlavacek-Larrondo et al., 2020; Barfety et al., 2022), in some cases predicting as much as a MW-worth of stars contributed to the CCM by such cooling events.
@@
Some observations of isolated HII regions in nearby clusters, such as the Virgo cluster, may also hint at a more sustained but lower-level star formation activity taking place directly in the CCM/ICL region (Gerhard et al., 2002; Bellazzini et al., 2018). However, it is not always trivial to disentangle whether these isolated star-forming clouds are CCM-born material or belong to some older ram-pressure or interaction events in the cluster (Junais et al., 2021; Jones et al., 2022).
@@
Numerical simulations therefore emerge as valuable tools to explore this mode of star formation within the realistic framework of the CDM model. For instance, Puchwein et al. (2010) estimate that approximately of the ICL stars formed in their clusters (with masses between and ) originated at distances consistent with being CCM-born.
@@
Various numerical methods have been employed to study the formation of cold gas in the CCM, which could give rise to star formation outside galaxies. Runaway local production of cold gas in the CCM, as well as thermal instability, have been shown to lead to multi-phase structures in idealized hot haloes (Maller & Bullock, 2004; Kaufmann et al., 2006, 2009; McCourt et al., 2012; Sharma et al., 2012).
@@
Given that the cold-phase CCM resulting from such processes is a small-scale phenomenon, it has been studied largely in idealized numerical experiments (Fielding et al., 2020). Previous work by Oppenheimer et al. (2018) demonstrated a relatively significant abundance of cold, low-ionization gas in the CCM of EAGLE galaxies, albeit with uncertainties regarding the numerical robustness of the hydrodynamical technique in this regime (see also Ford et al. 2013).
@@
Semi-analytic models have also been developed to explain the unique, possibly drag-induced sub-virial kinematics of cold gas flowing through luminous red galaxy host haloes (Afruni et al., 2019). In the context of IllustrisTNG, Nelson et al. (2018) showed that TNG is consistent with empirical OVI constraints in the CCM.
@@
The cold gas content of galaxies themselves (both neutral HI and molecular H2) has also been favorably compared against data in several regimes. Nelson et al. (2020) also demonstrated that group mass haloes hosting massive galaxies are filled with a hot, virialized plasma at K, and that they contain a large abundance of cold gas with a characteristic temperature of K. This cold phase of the CCM takes the form of thousands of small, kpc-sized clouds, which fill the inner halo out to hundreds of kpc (similar to predictions by Maller & Bullock 2004 for MW-mass halos).
@@
There are, however, several important caveats when studying the formation of star-forming clouds in the CCM, mostly connected with limitations of numerical techniques. Approaches such as SPH have been shown to be unable to track the evolution of instabilities, such as the Kelvin-Helmholtz instability, leading to the formation of more dense clouds that are artificially long-lived compared to similar numerical experiments run with grid-base techniques (Agertz et al., 2009; Torrey et al., 2012).
@@
The lack of, or inadequate treatment of other important physics, such as heat conduction, magnetic fields or metal diffusion might also complicate the interpretation of results from simulations (see the discussion in Puchwein et al. 2010). It is clear that, for any numerical technique chosen, high resolution becomes key to study in-situ star formation in low-mass clouds embedded in the hot CCM of clusters, especially considering the low masses of these clouds.
@@
In this context, we present here a study of the in-situ formation of stars and their contribution to the ICL using the three most massive groups/clusters in the TNG50 simulations (Pillepich et al., 2018b; Nelson et al., 2019), one of the highest-resolution cosmological simulations of low-mass clusters with virial mass at .
@@ 2 Simulation
We use the cosmological hydrodynamical simulation IllustrisTNG (Pillepich et al., 2018b; Springel et al., 2018; Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2019). The IllustrisTNG project111https://www.tng-project.org is a suite of state-of-the-art cosmological galaxy formation simulations that come in different physical sizes and mass resolutions. In this work we use the highest resolution baryonic run TNG50-1 (TNG50 for short), which has a volume of approximately . It has resolution elements, and the baryon and dark matter particle masses are and , respectively. The IllustrisTNG cosmological parameters are consistent with CDM model determined by Planck XIII (Planck Collaboration et al., 2016) to be with and
Gravity and baryons are followed using the Arepo moving mesh code (Springel, 2010). The baryonic treatment is based on the code used for its predecessor simulation suite, Illustris (Vogelsberger et al., 2013a, 2014a, 2014b; Genel et al., 2014; Nelson et al., 2015), with modifications implemented to better track the formation and evolution of galaxies (as described in Pillepich et al., 2018a).
The updated IllustrisTNG sub-grid models account for star formation, radiative metal cooling, chemical enrichment from SNII, SNIa, and AGB stars, stellar feedback, and super-massive black hole feedback. The star formation in the Interstellar Medium (ISM) is treated in a sub-resolution fashion following a variation of the Springel & Hernquist (2003) with the adoption of a Chabrier (2003) initial mass function, and model the star-forming dense ISM gas using an effective equation of state where stars form stochastically above a gas density of = with a star formation time-scale of = Gyr (refer to Vogelsberger et al. 2013b for detailed explanations). These models were calibrated to reproduce a set of observational constraints that include the observed galaxy stellar mass function (Vogelsberger et al., 2014a), the cosmic star formation rate (SFR) density (Genel et al., 2014), colors and morphologies of the present day galaxy population (Sales et al., 2014; Snyder et al., 2015; Rodriguez-Gomez et al., 2017).
@@ 2.1 Defining ICL and cleaning method
2.2 H predictions
In order to compare results with the recent observations of the ICL in nearby clusters (i.e. Virgo), we use the calibration of Kennicutt & Evans (2012) to convert the star formation rates to H luminosities. These calibrations are based on evolutionary population synthesis models, in which the emergent spectral energy distributions (SEDs) are derived for synthetic stellar populations with a prescribed age mix, chemical composition, and Kroupa (2001) initial mass function (IMF). The resulting relation is
@@ 3 Results
We begin by characterizing the mass distribution within our groups. Fig. 2 illustrates the projected radial distribution of stars (solid thick black line), dark matter (black dotted line), and satellite galaxies with (depicted as orange circles) after projecting each group along the axis. The stellar component is predicted to fall off more steeply than both the dark matter and the satellite galaxies in these groups (upon vertical re-scaling by a factor of , the satellite galaxies appear to closely trace the dark matter distribution in this simulation). For comparison, the purple dotted line represents the stellar mass surface density profiles from Cooper et al. (2013), averaged for halos with . These results align closely with our stellar profiles (solid black lines) for the two less massive clusters. However, in the case of our most massive cluster, having a halo mass twice that of the most massive halo in their sample, differences with the predictions in Cooper et al. (2013) are evident. Note that the tagging method in Cooper et al. (2013) does not properly account for the in-situ star formation in the BCG, therefore giving a flatter slope in the inner regions. We further divide the stars according to their origin: in-situ (represented by the thin blue dot-dashed curve) and accreted (depicted by the thin red long-dashed curve). While the dominance of the in-situ population is expected within the BCG (the region highlighted by gray shading), their relatively substantial contribution to the ICL at hundreds of kpc is intriguing. In our systems, in-situ born stars contribute between and of the stellar mass in the ICL (refer to the top right corner of each panel for the specific percentage of each group).
@@
An inspection of the birth place of the in-situ labelled stars in the ICL shows that the majority of them are formed already at large distances from the BCG (although not associated with substructures) and are not stars from the central galaxy kicked-off by mergers, a common formation path for in-situ stars in the diffuse stellar halos of MW-like galaxies, and as is the case for our most massive cluster (which experienced a recent major merger in its formation history). However, for the remaining groups in our sample, over 60% of the in-situ ICL is formed at kpc or (where distance is in physical kpc and is measured with respect to the BCG or its main progenitor at the time of birth of the star). This suggests that gas in the CCM must somehow become available for star formation beyond the inner regions dominated by the BCG. This star formation mode is thus happening in unconventionally diffuse conditions, very different from the star formation resulting from the dense interstellar medium of galaxies.
@@
We calculate the total SFR of CCM gas at for our groups and compare this to the level of star formation ongoing in the central BCG as well as in other galaxies. Fig. 3 highlights that, taken as a unit, the ICL (integrated within the virial radius) is forming stars at rates comparable to that of the BCG and other massive galaxies. For this comparison, we assign a stellar mass to the ICL equal to the sum of all in-situ and accreted stellar particles that lay beyond and up to the virial radius and that are not associated to any satellite or identified substructure by SUBFIND. Similarly, we compute its SFR by summing the star formation rates of all gas cells that are not gravitationally bound to any substructure within the same region. Fig. 3 shows that while the ICL lays systematically below the “main sequence” of star forming galaxies (magenta points show central galaxies within the TNG50 box), its SFR is comparable to many of the massive satellite galaxies in the group (gray circles) and slightly above the SFR of the central BCG in the case of Group 1 (middle panel). For comparison, we present the extrapolated “main-sequence” relation derived from the galaxies in the Virgo cluster (see Boselli et al. 2023), for a sample with normal HI gas content (HI-normal) and deficient HI gas content (HI-deficient); the mean values and standard deviations for HI-normal and HI-deficient objects are shown by blue and red markers. We are also presenting results from late-type nearby galaxies in the Herschel Reference Survey by Boselli et al. 2015 (dashed green line), a large statistical sample of SDSS galaxies from Peng et al. 2010 (dashed brown line), and the extrapolated time-dependent main-sequence relation derived by Speagle et al. 2014 (dashed orange line).
@@
The ongoing star formation in the CCM shows a decreasing trend with distance from the center of the cluster. Fig. 4 presents the surface density of star formation rate (), computed by projecting our groups along the -axis and integrating star formation within square bins of kpc on each side. Different colors represent different groups, while individual points of the same color depict the scatter between various regions within a given group. Bins with no star formation at all are not shown. The thick solid curve corresponds to the median trend (considering all groups collectively) of as a function of clustercentric distance, highlighting that the inner regions of the CCM exhibit more prominent star formation compared to the outskirts.
@@
To ease comparison with observations, we convert into a surface density of H luminosity (right vertical axis) as explained in detail in Sec. 2.2 and following Kennicutt & Evans (2012). Typical values in the CCM of our sample span with considerable scatter. These values are lower than the average measured in central regions of disk galaxies, but comparable to the values in the outskirts of MW- and LMC-like objects (Tacchella et al., 2022; Bundy et al., 2015). Beyond clustercentric distances diffuse star formation is rare in all our systems.
@@
Projected maps of our halos in Fig. 5 show in more detail where the star formation in the intracluster space is happening. Panels on the first row show the star-forming gas in the region (bounds shown with magenta and orange circles) color coded by SFR (color map on the right). Star formation at is widespread in these systems and follows filament-like patterns. We highlight that special care was taken to remove any material gravitationally associated with surviving satellites, as described in Sec. 2.1, and this star formation is occurring in gas seemingly associated to the central host halo and not to bound substructures. In fact, further inspection of each system revealed that the filamentary nature of the star-forming gas follows the distribution of neutral gas in the main host halo (bottom row), and looks loosely linked to the dark matter distribution (middle row) for at least 2 of the systems (groups 1 and 2).
@@
To gain further insight, we zoom into smaller regions of group 1 in Fig. 6. The yellow points show the distribution of the stellar particles in each box and reveal the presence of several galaxies, some with associated stellar streams—most prominently observable in the two innermost regions (red and green panels). The presence of the ICL is apparent in panels as a diffuse yellow stellar background on the end closer to the central region of cluster. We highlight the presence of star forming gas colored in magenta by their estimated flux (refer to the color bar on the side of the main panel). For the zoomed-in panels, the color maps are calculated in kpc bins to align with the arcsec spatial resolution of the VESTIGE survey of the Virgo cluster (Boselli et al., 2018a), while the main panel is smoothed to 1 arcmin.
@@
The zoomed-in regions in Fig. 6 show that the majority of the star forming gas is mostly unrelated to the distribution of stars or galaxies. Star forming regions are also not obviously associated to ram-pressure tails from gas-rich galaxies note that this is different from what has been observed in cases such as Smith et al. (2010); Mihos et al. (2016); Boselli et al. (2018a); Gullieuszik et al. (2020), where the association of the star-forming region to tails of ram-pressure stripped galaxies is clear upon visual inspection. Instead, stars seem to be forming in cloudlet-like objects with tens-of-pc typical size, which condensate along filamentary structures in the CCM.
@@
This is not the first time that small-scale cold clouds in the CCM have been reported in TNG50. In a detailed analysis of the CCM in halos, Nelson et al. (2020) highlighted this fundamental prediction from the TNG model in combination with the high resolution achieved by TNG50. These small-scale cold clouds seem consistent with being seeded by local inhomogeneities in gas density that later grow due to increased cooling. While the role of magnetic fields in forming these clouds is not clear, the magnetic pressure in these clouds is large and exceeds the thermal pressure, providing support and making these features long-lived, on the scale of several Gyrs. Interestingly, the results in this work show that some of these cold clouds manage to become star-forming and contribute a non-negligible fraction of the stars to the diffuse ICL component.
@@
In agreement with the findings of Nelson et al. (2020), the metallicity of in-situ born ICL stars (with ) is relatively high, registering a median value of with a substantial scatter of approximately dex, comparable to the accreted ICL, the metallicity of dwarf galaxies with , and the outskirts of more luminous, MW-like galaxies. This suggests that the gas seeding these density perturbations has galaxy-like abundances and it likely comprises gas that has, a long time ago, been stripped from infalling galaxies. Obvious associations to those satellites do not seem to persist until in our simulations, suggesting that the timescales for the clouds to form and become star-forming are comparable to the orbital times of the satellites.
@@
The physical conditions for the star-forming clouds seem comparable to those of normal star-forming regions in the ISM of galaxies. The median density of the star-forming cells flagged as part of the ICL is , which is just above the density threshold for star formation in the TNG baryonic treatment . Interestingly, the fraction of stellar mass born in-situ in the ICL seems quite independent of numerical resolution, as clusters selected in TNG100 show similar levels of star formation rate in the ICL (see Appendix A). The presence of these in-situ born stars in the ICL is therefore not driven by the increase in resolution from TNG50 to TNG100 but is instead a more fundamental consequence (or prediction) of the model.

### v2

@@ Abstract
We study the fraction of the intra-cluster light (ICL) formed in-situ in the three most massive clusters of the TNG50 simulation, with virial masses . We find that a significant fraction of ICL stars (-) are born in-situ. This amounts to a total stellar mass comparable to the central galaxy itself. Contrary to simple expectations, only a sub-dominant fraction of these in-situ ICL stars are born in the central regions and later re-distributed to more energetic orbits during mergers. Instead, many in-situ ICL stars form directly hundreds of kiloparsecs away from the central galaxy, in clouds condensing out of the circum-cluster medium. The simulations predict a present-date diffuse star formation rate of 1 /yr, with higher rates at higher redshifts. The diffuse star forming component of the ICL is filamentary in nature, extends for hundreds of kiloparsecs and traces the distribution of neutral gas in the cluster host halo. We discuss briefly how numerical details of the baryonic treatment in the simulation, in particular the density threshold for star formation and the equation of state, may play a role in this result. We conclude that a sensitivity of erg s-1 cm-2 arcsec-2 in Hα flux (beyond current observational capabilities) would be necessary to detect this diffuse star-forming component in galaxy clusters.
@@
The simplest models of galaxy formation suggest that most stars in a cluster should form in the high-density regions of individual galaxies. These models imply that the majority of stars in the intra-cluster light (ICL) originate from the tidal removal of stars from luminous galaxies with mass similar to the Milky Way (MW) (Puchwein et al., 2010; Cui et al., 2014; Cooper et al., 2015a; Contini et al., 2013, 2018, 2019; Montenegro-Taborda et al., 2023), with relatively small contributions from dwarf mass galaxies, particularly in group environments with halo virial masses (Ahvazi et al., 2024). Nevertheless, observational studies (Barfety et al., 2022; Hlavacek-Larrondo et al., 2020) and some numerical simulations (Puchwein et al., 2010) indicate that a non-negligible fraction of ICL stars are formed in-situ.
In-situ ICL stars could have formed in the brightest cluster galaxy (BCG) galaxy and been later pushed onto highly energetic orbits by mergers. An alternative, more intriguing, possibility is that stars form directly in the intra-cluster medium (ICM), far away from the high density regions of the BCG. This low-density mode of star formation is potentially interesting, as it might occur under very different physical conditions than the normal star formation in galaxies, with consequences for the multi-phase nature of the ICM and its metallicity, as well as for the age distribution of ICL stars, especially if in-situ star formation continues to the present day. From an observational perspective, if a substantial amount of stars form outside galaxies, this would have strong implications for wide-field surveys in the UV or Hα, which are especially sensitive to the presence of young stars.
@@
In-situ star formation in the ICM is also predicted to be widespread at high-redshift, where AGN feedback might be less powerful and unable to offset the short timescales for cooling in the dense regions of the ICM (Webb et al., 2015b). In such a scenario, spontaneous run-away cooling in the ICM might lead to an increase of density high enough to trigger violent star-formation events in locations outside galaxies (Hlavacek-Larrondo et al., 2020).
@@
Interestingly, observations of off-center star formation at large distances from the BCG are consistent with this scenario for a few high- clusters (Webb et al., 2015a; Hlavacek-Larrondo et al., 2020; Barfety et al., 2022), in some cases predicting as much as a MW-worth of stars contributed to the ICM by such cooling events.
@@
Some observations of isolated HII regions in nearby clusters, such as the Virgo cluster, may also hint at a more sustained but lower-level star formation activity taking place directly in the ICM/ICL region (Gerhard et al., 2002; Bellazzini et al., 2018). However, it is not always trivial to disentangle whether these isolated star-forming clouds are ICM-born material or belong to some older ram-pressure or interaction events in the cluster (Junais et al., 2021; Jones et al., 2022).
@@
Numerical simulations therefore emerge as valuable tools to explore this mode of star formation within the realistic framework of the CDM model. For instance, Puchwein et al. (2010) estimate that approximately of the ICL stars formed in their clusters (with masses between and ) originated at distances consistent with being ICM-born. The presence of in-situ formed stars in extended stellar halos is also documented extensively in the literature on the scale of MW-mass halos (Zolotov et al., 2009; Font et al., 2011; Tissera et al., 2012; Cooper et al., 2015b; Pillepich et al., 2015).
@@
Various numerical methods have been employed to study the formation of cold gas in the ICM, which could give rise to star formation outside galaxies. Runaway local production of cold gas in the ICM, as well as thermal instability, have been shown to lead to multi-phase structures in idealized hot haloes (Maller & Bullock, 2004; Kaufmann et al., 2006, 2009; McCourt et al., 2012; Sharma et al., 2012).
@@
Given that the cold-phase ICM resulting from such processes is a small-scale phenomenon, it has been studied largely in idealized numerical experiments (Fielding et al., 2020). Previous work by Oppenheimer et al. (2018) demonstrated a relatively significant abundance of cold, low-ionization gas in the circum-galactic medium (CGM) of EAGLE galaxies (for MW and group mass halos), albeit with uncertainties regarding the numerical robustness of the hydrodynamical technique in this regime (see also Ford et al. 2013). Similarly, Butsky et al. (2019) demonstrated evidence for the presence of cold gas in the outskirts of the galaxy clusters (in the ICM) using the high-resolution hydrodynamical galaxy cluster simulation, RomulusC.
@@
Semi-analytic models have also been developed to explain the unique, possibly drag-induced sub-virial kinematics of cold gas flowing through luminous red galaxy host haloes (Afruni et al., 2019). In the context of IllustrisTNG, Nelson et al. (2018b) showed that TNG is consistent with empirical OVI constraints in the CGM of MW and group mass halos.
@@
The cold gas content of galaxies themselves (both neutral HI and molecular H2) has also been favorably compared against data in several regimes, although simplifying assumptions are needed to derive these cold-phase fractions which are not self-consistently followed in typical galaxy-scale simulations. Nelson et al. (2020) also demonstrated that group mass haloes hosting massive galaxies are filled with a hot, virialized plasma at K, and that they contain a large abundance of cold gas with a characteristic temperature of K. This cold phase of the ICM takes the form of thousands of small, kpc-sized clouds, which fill the inner halo out to hundreds of kpc (similar to predictions by Maller & Bullock 2004 for MW-mass halos).
@@
There are, however, several important caveats when studying the formation of star-forming clouds in the ICM, mostly connected with limitations of numerical techniques. Approaches such as SPH have been shown to be unable to track the evolution of instabilities, such as the Kelvin-Helmholtz instability, leading to the formation of more dense clouds that are artificially long-lived compared to similar numerical experiments run with grid-base techniques (Agertz et al., 2009; Torrey et al., 2012).
@@
The lack of, or inadequate treatment of other important physics, such as heat conduction, magnetic fields or metal diffusion might also complicate the interpretation of results from simulations (see the discussion in Puchwein et al. 2010). It is clear that, for any numerical technique chosen, high resolution becomes key to study in-situ star formation in low-mass clouds embedded in the hot ICM of clusters, especially considering the low masses of these clouds.
@@
In this context, we present here a study of the in-situ formation of stars and their contribution to the ICL using the three most massive groups/clusters in the TNG50 simulations (Pillepich et al., 2019; Nelson et al., 2019b), one of the highest-resolution cosmological simulations of low-mass clusters with virial mass at .
@@ 2 Simulation
We use the cosmological hydrodynamical simulation IllustrisTNG (Pillepich et al., 2018b; Springel et al., 2018a; Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2018a, 2019a). The IllustrisTNG project1111https://www.tng-project.org is a suite of state-of-the-art cosmological galaxy formation simulations that come in different physical sizes and mass resolutions. In this work we use the highest resolution baryonic run TNG50-1 (TNG50 for short) (Pillepich et al., 2019; Nelson et al., 2019b), which has a volume of approximately . It has resolution elements, and the baryon and dark matter particle masses are and , respectively. The IllustrisTNG cosmological parameters are consistent with CDM model determined by Planck XIII (Planck Collaboration et al., 2016) to be with and
Gravity and baryons are followed using the Arepo moving mesh code (Springel, 2010). The baryonic treatment is based on the code used for its predecessor simulation suite, Illustris (Vogelsberger et al., 2013a, 2014a, 2014b; Genel et al., 2014; Nelson et al., 2015), with modifications implemented to better track the formation and evolution of galaxies (as described in Pillepich et al., 2018a; Weinberger et al., 2017).
The updated IllustrisTNG sub-grid models account for star formation, radiative metal cooling, chemical enrichment from SNII, SNIa, and AGB stars, stellar feedback, and super-massive black hole feedback (Weinberger et al., 2017; Pillepich et al., 2018a). The star formation in the Interstellar Medium (ISM) is treated in a sub-resolution fashion following a variation of the Springel & Hernquist (2003) with the adoption of a Chabrier (2003) initial mass function, and model the star-forming dense ISM gas using an effective equation of state where stars form stochastically above a gas density of = with a star formation time-scale of = Gyr (refer to Vogelsberger et al. 2013b for detailed explanations). These models were calibrated to reproduce a set of observational constraints that include the observed galaxy stellar mass - size (Genel et al., 2018), the color bimodality (Nelson et al., 2018a), galaxy clustering (Springel et al., 2018b) and stellar mass function (Pillepich et al., 2018a).
@@ 2.1 Defining ICL and cleaning method
2.2 Hα predictions
In order to compare results with the recent observations of the ICL in nearby clusters (i.e. Virgo), we use the calibration of Kennicutt & Evans (2012) to convert the star formation rates to Hα luminosities. These calibrations are based on evolutionary population synthesis models, in which the emergent spectral energy distributions (SEDs) are derived for synthetic stellar populations with a prescribed age mix, chemical composition, and Kroupa (2001) initial mass function (IMF). The resulting relation is
@@ 3 Results
We begin by characterizing the mass distribution within our groups, focusing on the projected radial profiles to facilitate comparison with observational studies. Fig. 2 illustrates the projected radial distribution of stars (solid thick black line), and dark matter (black dotted line) after projecting each group along the axis. The stellar component is predicted to fall off more steeply than the dark matter component. For comparison, the purple dotted line represents the stellar mass surface density profiles from Cooper et al. (2013), based on a simplified particle tagging method, averaged for halos with . These results align closely with our stellar profiles (solid black lines) within the ICL region for the two less massive clusters (as shown in the right and middle panels). However, in the case of our most massive cluster, having a halo mass twice that of the most massive halo in their sample, differences with the predictions in Cooper et al. (2013) are expected (as a result the dashed purple line is removed from this panel). Note that the tagging method in Cooper et al. (2013) does not properly account for the in-situ star formation in the BCG, therefore giving a flatter slope in the inner regions. We further divide the stars according to their origin: in-situ (represented by the thin blue dot-dashed curve) and accreted (depicted by the thin red long-dashed curve). The relatively substantial contribution of the in-situ population to the ICL at hundreds of kpc is intriguing. In our systems, in-situ born stars contribute between and of the stellar mass in the ICL (refer to the top right corner of each panel for the specific percentage of each group).
@@
An inspection of the birth place of the in-situ labelled stars in the ICL shows that the majority of them are formed already at large distances from the BCG (although not associated with substructures) and are not stars from the central galaxy kicked-off by mergers, a common formation path for in-situ stars in the diffuse stellar halos of MW-like galaxies, and as is the case for our most massive cluster (which experienced a recent major merger in its formation history). However, for the remaining groups in our sample, over 60% of the in-situ ICL is formed at kpc or (where distance is in physical kpc and is measured with respect to the BCG or its main progenitor at the time of birth of the star). This suggests that gas in the ICM must somehow become available for star formation beyond the inner regions dominated by the BCG. This star formation mode is thus happening in unconventionally diffuse conditions, very different from the star formation resulting from the dense interstellar medium of galaxies.
@@
We calculate the total SFR of ICM gas at for our groups and compare this to the level of star formation ongoing in the central BCG as well as in other galaxies. Fig. 3 highlights that, taken as a unit, the ICL (integrated within the virial radius) is forming stars at rates comparable to that of the BCG and other massive galaxies. For this comparison, we assign a stellar mass to the ICL equal to the sum of all in-situ and accreted stellar particles that lie beyond and up to the virial radius and that are not associated to any satellite or identified substructure by SUBFIND. Similarly, we compute its SFR by summing the star formation rates of all gas cells that are not gravitationally bound to any substructure within the same region. Fig. 3 shows that while the ICL lies systematically below the extrapolation of the “main sequence” defined by lower-mass star forming galaxies (magenta points show central galaxies within the TNG50 box), its SFR is comparable to many of the massive satellite galaxies in the group (gray circles) and above the SFR of the central BCG in the case of Group 0 and Group 1 (left and middle panels). For comparison, we present the extrapolated “main-sequence” relation derived from the galaxies in the Virgo cluster (see Boselli et al. 2023), for a sample with normal HI gas content (HI-normal) and deficient HI gas content (HI-deficient); the mean values and standard deviations for HI-normal and HI-deficient objects are shown by blue and red markers. We are also presenting results from late-type nearby galaxies in the Herschel Reference Survey by Boselli et al. 2015 (dashed green line), a large statistical sample of SDSS galaxies from Peng et al. 2010 (dashed brown line), and the extrapolated time-dependent main-sequence relation derived by Speagle et al. 2014 (dashed orange line). One should also keep in mind that the fraction of quenched galaxies produced by the TNG model might overestimate the star formation in massive galaxies (e.g., see Fig. 8 in Donnari et al., 2021).
@@
The ongoing star formation in the ICM shows a decreasing trend with distance from the center of the cluster. Fig. 4 presents the surface density of star formation rate (), computed by projecting our groups along the -axis and integrating star formation within square bins of kpc on each side. Different colors represent different groups, while individual points of the same color depict the scatter between various regions within a given group. Bins with no star formation at all are not shown and not considered for calculating the median and the corresponding dipersion. The thick solid curve corresponds to the median trend (considering all groups collectively) of as a function of clustercentric distance, highlighting that the inner regions of the ICM exhibit more prominent star formation compared to the outskirts. We have checked that the vertical magenta structures in group 0 correspond to a few dominant clumps in the ICL with a wide range in predicted SFR.
@@
To ease comparison with observations, we convert into a surface density of Hα luminosity (right vertical axis) as explained in detail in Sec. 2.2 and following Kennicutt & Evans (2012). Typical values in the ICM of our sample span with considerable scatter. These values are lower than the average measured in central regions of disk galaxies, but comparable to the values in the outskirts of MW- and LMC-like objects (Tacchella et al., 2022; Bundy et al., 2015). Beyond clustercentric distances diffuse star formation is rare in all our systems.
@@
Projected maps of our halos in Fig. 5 show in more detail where the star formation in the intracluster space is happening. Panels on the first row show the star-forming gas in the region (bounds shown with magenta and orange circles) color coded by SFR (color map on the right). Star formation at is widespread in these systems and follows filament-like patterns. We highlight that special care was taken to remove any material gravitationally associated with surviving satellites, as described in Sec. 2.1, and this star formation is occurring in gas seemingly associated to the central host halo and not to bound substructures. In fact, further inspection of 3D projections for each system revealed that the filamentary nature of the star-forming gas follows the distribution of neutral gas in the main host halo (bottom row), and looks loosely linked to the dark matter distribution (middle row) for at least 2 of the systems (groups 1 and 2).
@@
To gain further insight, we zoom into smaller regions of group 1 in Fig. 6. The gray color map shows the distribution of the stellar component in each box and reveals the presence of several galaxies, some with associated stellar tails—most prominently observable in the two innermost regions (red and green panels). We highlight the presence of star forming gas colored in magenta by their estimated flux (refer to the color bar on the side of the main panel). For the zoomed-in panels, the color maps are calculated in kpc bins to align with the arcsec spatial resolution of the VESTIGE survey of the Virgo cluster (Boselli et al., 2018a), while the main panel is smoothed to 1 arcmin.
@@
The zoomed-in regions in Fig. 6 show that the majority of the star forming gas is mostly unrelated to the distribution of stars or galaxies. Star forming regions are also not obviously associated to ram-pressure tails from gas-rich galaxies. Note that this is different from what has been observed in cases such as Smith et al. (2010); Mihos et al. (2016); Boselli et al. (2018a); Gullieuszik et al. (2020), where the association of the star-forming region to tails of ram-pressure stripped galaxies is clear upon visual inspection. While these ram-pressure tails have been found in simulations including TNG50 (Rohr et al., 2023), they are not the main contributor to the star forming gas in our systems. Instead, stars seem to be forming in cloudlet-like objects with kpc typical size, which condensate along filamentary structures in the ICM.
@@
This is not the first time that small-scale cold clouds in the ICM have been reported in TNG50. In a detailed analysis of the ICM in halos, Nelson et al. (2020) highlighted this fundamental prediction from the TNG model in combination with the high resolution achieved by TNG50. These small-scale cold clouds seem consistent with being seeded by local inhomogeneities in gas density that later grow due to increased cooling. While the role of magnetic fields in forming these clouds is not clear, the magnetic pressure in these clouds is large and exceeds the thermal pressure, providing support and making these features long-lived, on the scale of several Gyrs. Interestingly, the results in this work show that some of these cold clouds manage to become star-forming and contribute a non-negligible fraction of the stars to the diffuse ICL component.
@@
In agreement with the findings of Nelson et al. (2020), the metallicity of in-situ born ICL stars (with ) is relatively high, registering a median value of with a substantial scatter of approximately dex, comparable to the accreted ICL, the metallicity of dwarf galaxies with , and the outskirts of more luminous, MW-like galaxies. This suggests that the gas seeding these density perturbations has galaxy-like abundances and it likely comprises gas that has, a long time ago, been stripped from infalling galaxies. Obvious associations to those satellites do not seem to persist until in our simulations, suggesting that the timescales for the clouds to form and become star-forming are comparable to the orbital times of the satellites. The similarity between the metallicities for the in-situ and accreted components complicates the chances of observationally distinguishing these two populations.
@@
The physical conditions for the star-forming clouds seem comparable to those of normal star-forming regions in the ISM of galaxies. The median density of the star-forming cells flagged as part of the ICM is , which is just above the density threshold for star formation in the TNG baryonic treatment (Please see Appendix A for a more detailed comparison of density and temperature for star-forming cells in the ICL compared to those in satellite galaxies.) Interestingly, the fraction of stellar mass born in-situ in the ICM seems only weakly dependent on numerical resolution, as some of the clusters selected in TNG100 show similar levels of star formation rate in the ICM (see Appendix B). The presence of these in-situ born stars in the ICL is therefore a fundamental consequence (or prediction) of the model in several resolution levels.